• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 18:43:57 on October 2,2009

    Entry number 292775

    keyword=Compton Status

    Alexandre increased the size of a buffer in the Compton counting DAQ and this allowed us to take two production runs where ROCs 1 and 2 crashed after one hour, instead of more quickly. He will increase the buffer size more to see if this fixes the problem for two hour runs.

    As soon as beam returns I will try to take some photon-electron coincidence data with the counting DAQ, and at the same time I will trigger the FADC with the electron detector. Right now spyconf and also the FADC flags file are set up to run in this test configuration.